Reports 1-1 of 1 Clear search Modify search
AdV-INJ (Optical setup for ITF reflection (SIB2))
chiummo, spinicelli, gouaty - 23:22 Tuesday 06 April 2021 (51336) Print this report
B2 photodiodes realignment
This evening Antonino and Piernicola adjusted the position of the mirror SIB2_M5 in order to remove some clipping (initially they observed a spot at the left corner of the mount with the infrared viewer). To this purpose they shifted the mirror to give more clearance for the transmitted beam.

After that we checked the centering of the three B2 photodiodes using the picomotors. B2_M5_H: +200 steps, B2_M6_H: +4000 steps, B2_M6_V:+4000 steps, B2_M5b_H:+2000 steps.

The attached figures compare the spectra of the audio channels in the three photodiodes. The B2_PD2 photodiode spectrum has been rescaled by a factor 0.2.
There is no obvious difference between the B2_PD3 spectrum and the other photodiodes any more which indicates that the clipping has been removed.
Images attached to this report
Search Help